# Generative Neural Video Compression via Video Diffusion Prior

### v1

@@ 4.2 Comparison Results
Quantitative Comparisons. We quantitatively evaluate GNVC-VD against representative learned and generative video codecs on UVG [UVG], MCL-JCV [MCL-JCV], and HEVC Class B [flynn16common]. As summarized in Fig. 4, GNVC-VD delivers consistent gains across perceptual quality metrics. On UVG, GNVC-VD achieves over 98% BD-rate reduction in DISTS and 56% in LPIPS compared with the distortion-oriented baseline DCVC-RT [DCVC-RT]. Compared with the generative codec GLC-Video [qi2025generative], GNVC-VD further reduces BD-rate by 86% in DISTS and 21% in LPIPS, as reported in Table 3. Beyond spatial perceptual quality, GNVC-VD also demonstrates superior temporal coherence, achieving higher CLIP-F scores and substantially lower $E_{\text{warp}}$ than GLC-Video, indicating more stable motion reconstruction and reduced frame-level flickering, as illustrated in Table 2.
@@ A.3 Neural-based Codecs
  • GLC-Video. We use the reconstructed videos and bitrates provided directly by the original authors of GLC-Video [qi2025generative]. All evaluation metrics are computed from the provided reconstructions.
  • PLVC. PLVC [plvc] is evaluated using its official implementation [^4] https://github.com/RenYang-home/PLVC and pre-trained weights. Since PLVC adopts HiFiC [HiFiC] for I-frame coding, we use its PyTorch implementation [^5] https://github.com/Justin-Tan/high-fidelity-generative-compression for consistency.
@@ Appendix B Model Implementation Details
Anchor latent (I-frame). The processing pipeline for the anchor latent is shown in Fig. 3(a). We adopt a design similar to ELIC [ELIC], where the analysis and synthesis transforms ($g_{s}$ and $g_{a}$) are constructed from cascaded residual bottleneck blocks [Resnet] and attention blocks [Cheng2020]. A joint space–channel context model estimates the probability distribution of the quantized anchor latent $\hat{y}_{1}$.
Predictive latents (P-frames). As illustrated in Fig. 3(b), for the predictive latents, we follow the architecture of DCVC-RT [DCVC-RT], where the transforms $g_{s}$ and $g_{a}$ are built from cascaded DC Blocks [DCVC-RT]. To balance coding efficiency and reconstruction quality, we adopt the two-step distribution estimation scheme described in [DCVC-HEM].
@@ C.1 Additional Metrics Evaluation
For a more comprehensive comparison, we report the rate–distortion curves of all baseline methods and our GNVC-VD in terms of PSNR, MS-SSIM, and LPIPS-Alex in Fig. 9. The VGG-based LPIPS variant correlates more strongly with human perception in generative compression. Therefore, in the main paper, perceptual comparisons are reported using LPIPS-VGG, which provides a more reliable indicator of perceptual fidelity. However, because the AlexNet-based LPIPS metric is more commonly used in the learned compression literature, we additionally include LPIPS-Alex results here for completeness. Compared with perceptual codecs such as GLC-Video [qi2025generative] and PLVC [plvc], GNVC-VD achieves clear improvements in distortion-oriented metrics (PSNR and MS-SSIM) while also delivering notably better perceptual quality (LPIPS-Alex), consistent with the LPIPS-VGG and DISTS improvements reported in the main paper. Relative to MSE-optimized codecs, although a small gap remains in PSNR and MS-SSIM, GNVC-VD provides substantially superior perceptual fidelity.

### v2

@@ 4.2 Comparison Results
Quantitative Comparisons. We quantitatively evaluate GNVC-VD against representative learned and generative video codecs on UVG [35], MCL-JCV [48], and HEVC Class B [8]. As summarized in Fig. 4, GNVC-VD delivers consistent gains across perceptual quality metrics. On UVG, GNVC-VD achieves over 98% BD-rate reduction in DISTS and 56% in LPIPS compared with the distortion-oriented baseline DCVC-RT [16]. Compared with the generative codec GLC-Video [38], GNVC-VD further reduces BD-rate by 86% in DISTS and 21% in LPIPS, as reported in Table 3. Beyond spatial perceptual quality, GNVC-VD also demonstrates superior temporal coherence, achieving higher CLIP-F scores and substantially lower $E_{\text{warp}}$ than GLC-Video, indicating more stable motion reconstruction and reduced frame-level flickering, as illustrated in Table 2.
@@ A.3 Neural-based Codecs
  • GLC-Video. We use the reconstructed videos and bitrates provided directly by the original authors of GLC-Video [38]. All evaluation metrics are computed from the provided reconstructions.
  • PLVC. PLVC [52] is evaluated using its official implementation [^4] https://github.com/RenYang-home/PLVC and pre-trained weights. Since PLVC adopts HiFiC [33] for I-frame coding, we use its PyTorch implementation [^5] https://github.com/Justin-Tan/high-fidelity-generative-compression for consistency.
@@ Appendix B Model Implementation Details
Anchor latent (I-frame). The processing pipeline for the anchor latent is shown in Fig. 3(a). We adopt a design similar to ELIC [11], where the analysis and synthesis transforms ($g_{s}$ and $g_{a}$) are constructed from cascaded residual bottleneck blocks [12] and attention blocks [5]. A joint space–channel context model estimates the probability distribution of the quantized anchor latent $\hat{y}_{1}$.
Predictive latents (P-frames). As illustrated in Fig. 3(b), for the predictive latents, we follow the architecture of DCVC-RT [16], where the transforms $g_{s}$ and $g_{a}$ are built from cascaded DC Blocks [16]. To balance coding efficiency and reconstruction quality, we adopt the two-step distribution estimation scheme described in [24].
@@ C.1 Additional Metrics Evaluation
For a more comprehensive comparison, we report the rate–distortion curves of all baseline methods and our GNVC-VD in terms of PSNR, MS-SSIM, and LPIPS-Alex in Fig. 9. The VGG-based LPIPS variant correlates more strongly with human perception in generative compression. Therefore, in the main paper, perceptual comparisons are reported using LPIPS-VGG, which provides a more reliable indicator of perceptual fidelity. However, because the AlexNet-based LPIPS metric is more commonly used in the learned compression literature, we additionally include LPIPS-Alex results here for completeness. Compared with perceptual codecs such as GLC-Video [38] and PLVC [52], GNVC-VD achieves clear improvements in distortion-oriented metrics (PSNR and MS-SSIM) while also delivering notably better perceptual quality (LPIPS-Alex), consistent with the LPIPS-VGG and DISTS improvements reported in the main paper. Relative to MSE-optimized codecs, although a small gap remains in PSNR and MS-SSIM, GNVC-VD provides substantially superior perceptual fidelity.